OpenStudy (anonymous):

2over3 + 3mover5, = 31over15. (These are fractions.) What is m?

OpenStudy (anonymous):

m=21/9

OpenStudy (anonymous):

(2/3) + (3m/5) = (31/15) => (10/15) + (9m/15) = (31/15) => 10 + 9m =31 => 9m = 21 => m = 21/9